# Evaluating and Steering Modality Preferences in Multimodal Large Language Model

### v2

@@ Abstract
Multimodal large language models (MLLMs) have achieved remarkable success on complex multimodal tasks. However, it remains insufficiently explored whether they exhibit modality preference, a tendency to favor one modality over another when processing multimodal contexts. To study this question, we introduce MC2 benchmark, which constructs controlled evidence-conflict scenarios to systematically evaluate modality preference in decision-making. Extensive experiments reveal that all 20 tested MLLMs generally demonstrate clear modality preferences, and such preferences can serve as a useful indicator of downstream task performances of MLLMs. Further analysis shows that modality preference can be controlled by instruction guidance and captured within the latent representations of MLLMs. Built on these insights, we propose a probing and steering method based on representation engineering to explicitly control modality preference without requiring additional fine-tuning. This method effectively amplifies modality preference toward a desired direction and demonstrates promising improvements across multiple downstream applications, including multimodal visual understanding and multimodal machine translation. footnotetext: The code and data are released at https://github.com/EchoDreamer/Modality-Preference.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs; Achiam et al., 2023; Team et al., 2023; Wang et al., 2024; Yin et al., 2024) have emerged as a powerful paradigm for processing and reasoning across heterogeneous data modalities (e.g., text, images, video). Recent advances demonstrate their exceptional capabilities on complex tasks with multimodal contexts, including autonomous web browsing (He et al., 2024), graphical user interface understanding (Hong et al., 2024b), and multimodal dialogue systems (Sun et al., 2022). Despite impressive performance, fundamental questions remain about their modality preference—whether MLLMs tend to rely more heavily on one modality than others, and to what extent they favor a specific modality when resolving multimodal inputs.
To investigate this, one line of work (Fu et al., 2024; Amara et al., 2024) compares model performance on unimodal input, providing either only text or only image input for the same question. Another line of research analyzes the relative contributions of textual and visual context, typically by removing one modality to observe the changes of the downstream performance (Park et al., 2025) or Shapley value (Alishahi et al., 2019; Parcalabescu & Frank, 2024; 2022). However, such settings inherently introduce bias, as they isolate modalities, thus failing to reflect how models process inputs in realistic multimodal scenarios, where information from different modalities naturally co-occur.
In this paper, we provide a controllable setup to study the modality preference in MLLMs. As shown in the left panel of Figure 1, we introduce a modality context conflict setting, where MLLMs are asked to answer a question based on a pair of contrasting evidence from different modalities. In this way, we can determine the modality preference based on the answer given by MLLMs. To dismiss the effect of confounding factors including question comprehension, single-modality perception, and the internal knowledge of MLLMs, we annotate and select perception-level tasks that demonstrate accurate question comprehension and reliable single-modality recognition. Building upon this, we introduce a semi-automated annotation framework to construct a refined Modality Context Conflict dataset, MC2, which covers eight perception-level tasks with 2,000 carefully selected samples. Using MC2, we conduct a comprehensive analysis of modality preference across a diverse set of 20 representative MLLMs. Our study reveals several intriguing findings:
  • Most MLLMs (except Qwen2.5-VL and InternVL3) display text preference, as shown in the right panel of Figure 1, and modality preference can serve as a useful indicator of downstream task performances.
@@
Built on these, we propose a modality preference probing and steering method based on representation engineering (Zou et al., 2023) to explicitly amplify the modality preference without additional fine-tuning. Experimental results show that the proposed method leads to notable performance improvements on multimodal visual understanding and multimodal machine translation. Our main contributions are summarized as follows:
Figure 1: Illustrations of evaluating modality preference. Left: Using multimodal conflict context pairs to evaluate modality preference. Right: Quantified scores for Vision and Text modalities, where a higher score indicates a stronger preference toward the corresponding modality. Other represents irrelevant predictions, which are discarded during the statistics.
@@
  • Our analysis reveals that intrinsic modality preferences in MLLMs are steerable and identifiable through latent representation, providing insights into multimodal reasoning.
@@ 2.1 Modality Preference
Data factors influencing modality preference. Research on data-related factors (Guo et al., 2023; Chen et al., 2024; Leng et al., 2024) explores how properties of multimodal datasets give rise to and reinforce modality preference. In particular, many samples in multimodal datasets can be resolved correctly by relying on information from only a single modality. When prevalent in training data, such samples bias the optimization dynamics, encouraging models to disproportionately rely on a single modality (Chen et al., 2024). Furthermore, their inclusion in evaluation benchmarks artificially inflates performance metrics, as models can exploit these unimodal shortcuts instead of performing genuine cross-modal integration (Leng et al., 2024; Winterbottom et al., 2020). While these studies establish data’s role in inducing preference, our work focuses on exploring the intrinsic modality preference inherent in MLLMs themselves, independent of specific data distributions.
Evaluating the intrinsic modality preference in MLLMs. Early studies (Peng et al., 2022; Yang et al., 2024; Wei et al., 2024) analyze modality preference or bias by examining how multimodal models optimize for multimodal inputs. Through such analyses, researchers observe that modality bias has a significant impact on both model optimization and downstream task performance (Peng et al., 2022; Ren et al., 2022; Zhang et al., 2024). While these works offer valuable insights, they typically require training models from scratch, which makes them impractical for large-scale multimodal systems. Recent studies have investigated intrinsic modality preference in MLLMs by evaluating model performance on unimodal inputs—using only text or only image for the same task (Fu et al., 2024; Amara et al., 2024)—and by applying Shapley value-based attribution methods to quantify the contribution of each modality (Alishahi et al., 2019; Parcalabescu & Frank, 2022; 2024). However, in real-world multimodal applications, all modalities are indispensable for task resolution, making these frameworks inadequate for determining truly modality preference. Wu et al. (2025) evaluate the model’s ability to detect conflict under scenarios involving conflicting multimodal contexts. However, conflict detection is only one facet of multimodal reasoning and does not comprehensively reflect a model’s modality preference when processing multimodal contexts.
In this work, we simulate multimodal reasoning by examining the behavior of MLLMs in response to questions under scenarios involving conflicting multimodal contexts. Compared to prior work, we carefully control confounding variables such as input quality, question-understanding ability, and internal model knowledge, and construct a modality context conflict dataset, enabling a more rigorous evaluation of modality preference and uncovering new insights. Furthermore, we design a flexible method which can controllably steer the modality preference and demonstrate effectiveness across multiple downstream tasks.
@@ 2.2 Representation Engineering
Extensive research has shown that large language models (LLMs) encode interpretable concepts, such as sentiment, truthfulness, and stylistic attributes in representation space in LLMs (Liu et al., 2023; Panickssery et al., 2023; Subramani et al., 2022; Turner et al., 2023). Building on this foundation, representation engineering has proven effective for editing, enhancing, or suppressing specific behaviors in LLMs (Greenblatt et al., 2023; Stolfo et al., 2024; Wu et al., 2024; Xu et al., 2024; Zou et al., 2023). In this work, we extend this paradigm to a novel setting: controlling modality preference in multimodal large language models (MLLMs). Instead of focusing on abstract properties, our method identifies and manipulates representation directions that are sensitive to modality preference, enabling flexible and targeted control over multimodal reasoning behavior.
@@ 3.1 Data Design Philosophy
To reveal the modality preference of MLLMs when processing truly multimodal inputs, we construct modality context conflict pairs and carefully control various confounding factors that may affect the measurement of modality preference, such as the internal knowledge of MLLMs, question comprehension, and the understanding of unimodal contexts. Therefore, we focus on evaluating perception-level tasks that require minimal external knowledge, reducing the impact of internal knowledge within MLLMs. Besides, we ensure that the evaluated models can accurately answer questions based solely on the context from a single modality, minimizing the influence of confounding factors.
@@ 3.2 Semi-automated Data Construction Pipeline
In this section, we introduce our semi-automated data construction pipeline, which follows a meticulous and iterative process to ensure the robustness and reliability of the dataset, in line with the design philosophy outlined in Section 3.1. The dataset is derived from the TDIUC (Kafle & Kanan, 2017) dataset, sourced from MS-COCO (Lin et al., 2014), widely adopted in model development to ensure the evaluated models can recognize the images. We select the image as vision context $c^{v}$, question $q$, and answer $A^{v}$ based on the vision context and the image caption capcap for each sample from TDIUC as the foundation for data annotation. The pipeline follows these steps:
Textual Context Construction. Given a sample including $c^{v}$, $q$, $A^{v}$ and capcap, we construct candidate contrastive textual contexts $c^{t}$ that conflict with $c^{v}$ specifically in relation to $q$ but are aligned with the $c^{v}$ and capcap in terms of overall scene semantics. We prompt DeepSeekV3 (Liu et al., 2024a) and ChatGPT4o-mini (Hurst et al., 2024b) to generate a distractor answer $A^{t}$ to $q$, together with $c^{t}$ that plausibly supports $A^{t}$, using carefully crafted instructions.
For each model, we generate two pairs of $A^{t}$ and $c^{t}$ to facilitate downstream data selection. To ensure that all evaluated MLLMs demonstrate strong recognition capabilities for both visual and textual contexts, we employ several basic MLLMs, such as LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023), as judges to select samples that can be correctly understood with respect to $c^{v}$ and $c^{t}$.
@@
Iterative Refinement. The dataset undergoes multiple rounds of refinement through a feedback loop between textual context generation and human verification, which helps identify and rectify potential errors, thereby enhancing the dataset quality.
Modality Context Conflict Dataset. To this end, we construct MC2, a modality context conflict dataset including 2000 samples.
The detailed instruction templates for textual context generation, the detailed manual annotation procedures, the data annotation format along with sample cases and dataset statistics are provided in Appendix B.
@@ 3.3 Question Design and Evaluation Metric
Question Design. We reformulate the original questions with ChatGPT-4o-mini (Hurst et al., 2024a) into a binary-choice format. To further reduce potential position bias in multimodal inputs, we adopt a consistent evaluation strategy, similar to Liu et al. (2024d). Concretely, for each question, we construct two versions by swapping the order of the answer choices. A model’s prediction is regarded as consistent only if it selects the same answer in both versions for a sample; otherwise, it is labeled as inconsistent. Such inconsistent samples are discarded from the subsequent measurement of modality preference.
Evaluation Metric. Inspired by prior work on evaluating stylistic or knowledge-related preferences of LLMs and MLLMs through conflict-pair contexts (Li et al., 2024b; Xie et al., 2023; Liu et al., 2025), we extend this idea to evaluate the modality preference by designing a metric that captures how MLLMs respond to conflicting signals from different modalities. More importantly, through the careful design of our benchmark, we establish as a basis that the model can reliably understand both modalities in isolation. As shown in Table 9 and Table 10 in Appendix, all models achieve over 95% accuracy when provided with either textual or visual context.
Building on this, our metric evaluates modality preference by assessing how the model’s responses align with textual or visual input when the two provide conflicting signals. The model’s response is then categorized based on which modality it aligns with: 1) Vision: the response aligns with the visual context; 2) Text: the response aligns with the textual context; 3) Others: the responses are ambiguous, uncertain, or inconsistent with either modality, which are discarded from further analysis. Then, we naturally define the Vision Ratio to quantify the model’s preference toward the vision modality, defined as: ${S_{\text{vision}}}/({S_{\text{vision}}+S_{\text{text}}})$, where $S_{\text{vision}}$ or $S_{\text{text}}$ denotes the score of the vision or text modality, computed as the proportion of samples whose responses are categorized as Vision or Text across the dataset. $S_{others}$ is the proportion of samples whose responses are categorized as Others. Vision Ratio greater than 0.5 indicates that the model tends to favor visual context over text.
@@ 4 Modality Preferences in MLLMs
This section presents a systematic investigation into modality preference in MLLMs, structured around four key research questions: 1) Which modality do MLLMs prioritize? 2) What factors drive these preferences? 3) Can the Vision Ratio provide guidance for downstream task performance? 4) Can modality preference be controlled? This investigation helps uncover the underlying mechanisms of modality preference and enables us to apply these insights to downstream tasks.
Figure 2: Results of modality preference across different MLLMs. Left: Quantified scores for Vision, Others Text modalities using $S_{vision}$, $S_{others}$ and $S_{text}$ as well as Vision Ratio. Right: Trends of Vision Ratio with respect to model parameter size across different MLLMs.

### v3

# Evaluating and Steering Modality Preferences in Multi-modal LLMs

@@ Abstract
Multi-modal large language models (MLLMs) have achieved remarkable success on complex multi-modal tasks. However, it remains insufficiently explored whether they exhibit modality preference, a tendency to favor one modality over another when processing multi-modal contexts. To study this question, we introduce MC2 benchmark, which constructs controlled evidence-conflict scenarios to systematically evaluate modality preference in decision-making. Extensive experiments reveal that all 20 tested MLLMs generally demonstrate clear modality preferences, and such preferences can serve as a useful indicator of downstream task performance of MLLMs. Further analysis shows that modality preference can be controlled by instruction guidance and captured within the latent representations of MLLMs. Built on these insights, we propose a probing and steering method based on representation engineering to explicitly control modality preference without requiring additional fine-tuning. This method effectively amplifies modality preference toward a desired direction and demonstrates promising improvements across multiple multi-modal understanding and reasoning tasks. footnotetext: The code and data are released at https://github.com/EchoDreamer/Modality-Preference.
Machine Learning, ICML
@@ 1 Introduction
multi-modal Large Language Models (MLLMs; Achiam et al., 2023; Team et al., 2023; Wang et al., 2024; Yin et al., 2024) have emerged as a powerful paradigm for processing and reasoning across heterogeneous data modalities (e.g., text, images, video). Recent advances demonstrate their exceptional capabilities on complex tasks with multi-modal contexts, including autonomous web browsing (He et al., 2024), graphical user interface understanding (Hong et al., 2024b), and multi-modal dialogue systems (Sun et al., 2022). Despite impressive performance, fundamental questions remain about their modality preference—whether MLLMs tend to rely more heavily on one modality than others, and to what extent they favor a specific modality when resolving multi-modal inputs.
Figure 1: Illustrations of evaluating modality preference. Using multi-modal conflict context pairs to evaluate modality preference.
To investigate this, one line of work (Fu et al., 2024; Amara et al., 2024) compares model performance on unimodal input, providing either only text or only image input for the same question. Another line of research analyzes the relative contributions of textual and visual context, typically by removing one modality to observe the changes in the downstream performance (Park et al., 2025) or Shapley value (Alishahi et al., 2019; Parcalabescu and Frank, 2024, 2022). However, such settings inherently introduce bias, as they isolate modalities, thus failing to reflect how models process inputs in realistic multi-modal scenarios, where information from different modalities naturally co-occurs.
In this paper, we provide a controllable setup to study the modality preference in MLLMs. As shown in the Figure 1, we introduce a modality context conflict setting, where MLLMs are asked to answer a question based on a pair of contrasting pieces of evidence from different modalities. In this way, we can determine modality preference based on the answer given by MLLMs.
To enable a rigorous and fair assessment, we eliminate existing confounding factors including the internal knowledge of MLLMs and inherent understanding capabilities. Consequently, we curate and annotate the dataset to ensure that the model achieves precise question and single-modality context comprehension. Building upon this, we introduce a semi-automated annotation framework to construct a refined Modality Context Conflict dataset, MC2, which covers eight tasks with 2,000 carefully selected samples. Using MC2, we conduct a comprehensive analysis of modality preference across a diverse set of 20 representative MLLMs. Our study reveals several intriguing findings:
  • Most MLLMs (except Qwen2.5-VL and InternVL3) display text preference, as shown in the Figure 2, and modality preference can serve as a useful indicator of downstream task performance.
@@
Built on these, we propose a modality preference probing and steering method based on representation engineering (Zou et al., 2023) to explicitly amplify the modality preference without additional fine-tuning. Experimental results show that the proposed method leads to notable performance improvements on multi-modal understanding and reasoning tasks. Our main contributions are summarized as follows:
@@
  • Our analysis reveals that intrinsic modality preferences in MLLMs are steerable and identifiable through latent representations, providing insights into multi-modal reasoning.
@@ 2 Related Work
2.1 multi-modal Large Language Models
In recent years, Large Language Models (LLMs) have achieved impressive results across a wide spectrum of tasks (Lu et al., 2022; Zhang et al., 2024b; Guo et al., 2024; Xu et al., 2025; Huang and Chang, 2023). Building upon these, multi-modal Large Language Models (MLLMs) have extended such capabilities, demonstrating strong performance in visual question answering, structured information understanding, and visual navigation (Zhu et al., 2024; Zhang et al., 2025a; Liu and Liu, 2025; Wei et al., 2023; Zhang et al., 2025b). Recent research has turned to systematically evaluating their biases and failure modes (J. Liu, Y. Fu, R. Xie, R. Xie, X. Sun, F. Lian, Z. Kang, and X. Li (2024c); 11; K. Amara, L. Klein, C. Lüth, P. Jäger, H. Strobelt, and M. El-Assady (2024); A. Alishahi, G. Chrupała, and T. Linzen (2019); M. Chen, Y. Cao, Y. Zhang, and C. Lu (2024b)), aiming to better understand their limitations and guide the development of more reliable models.
@@ 2.2 Modality Preference
Data factors influencing modality preference. Research on data-related factors (Guo et al., 2023; Chen et al., 2024a; Leng et al., 2024) explores how properties of multi-modal datasets give rise to and reinforce modality preference. In particular, many samples in multi-modal datasets can be resolved correctly by relying on information from only a single modality. When prevalent in training data, such samples bias the optimization dynamics, encouraging models to disproportionately rely on a single modality (Chen et al., 2024a). Furthermore, their inclusion in evaluation benchmarks artificially inflates performance metrics, as models can exploit these unimodal shortcuts instead of performing genuine cross-modal integration (Leng et al., 2024; Winterbottom et al., 2020). While these studies establish data’s role in inducing preference, our work focuses on exploring the intrinsic modality preference inherent in MLLMs themselves, independent of specific data distributions.
Evaluating the intrinsic modality preference in MLLMs. Early studies (Peng et al., 2022; Yang et al., 2024; Wei et al., 2024) analyze modality preference or bias by examining how multi-modal models optimize for multi-modal inputs. Through such analyses, researchers observe that modality bias has a significant impact on both model optimization and downstream task performance (Peng et al., 2022; Ren et al., 2022; Zhang et al., 2024a). While these works offer valuable insights, they typically require training models from scratch, which makes them impractical for large-scale multi-modal systems. Recent studies have investigated intrinsic modality preference in MLLMs by evaluating model performance on unimodal inputs—using only text or only image for the same task (Fu et al., 2024; Amara et al., 2024)—and by applying Shapley value-based attribution methods to quantify the contribution of each modality (Alishahi et al., 2019; Parcalabescu and Frank, 2022, 2024). However, in real-world multi-modal applications, all modalities are indispensable for task resolution, making these frameworks inadequate for determining true modality preference. (Wu et al., 2025) evaluate the model’s ability to detect conflict under scenarios involving conflicting multi-modal contexts. However, conflict detection is only one facet of multi-modal reasoning and does not comprehensively reflect a model’s modality preference when processing multi-modal contexts.
In this work, we investigate the behavioral tendencies of MLLMs by examining their responses to questions within scenarios involving conflicting multi-modal contexts. Unlike prior studies, we decouple modality preference from confounding variables—such as internal model knowledge and varying understanding capabilities—by ensuring the models can accurately comprehend the provided context. Furthermore, we design a flexible method to steer the modality preference and demonstrate effectiveness across multiple downstream tasks.
@@ 2.3 Representation Engineering
Extensive research has shown that large language models (LLMs) encode interpretable concepts, such as sentiment, truthfulness, and stylistic attributes in representation space in LLMs (Liu et al., 2023b; Panickssery et al., 2023; Subramani et al., 2022; Turner et al., 2023). Building on this foundation, representation engineering has proven effective for editing, enhancing, or suppressing specific behaviors in LLMs (Greenblatt et al., 2023; Stolfo et al., 2024; Wu et al., 2024; Xu et al., 2024; Zou et al., 2023). In this work, we extend this paradigm to a novel setting: controlling modality preference in multi-modal large language models (MLLMs). Instead of focusing on abstract properties, our method identifies and manipulates representation directions that are sensitive to modality preference, allowing for flexible and precise modulation of model behavior.
@@ 3.1 Data Design Philosophy
Modality preference is a fundamental behavioral tendency to favor one modality over another, irrespective of MLLMs’s internal knowledge and inherent understanding capabilities. To enable a rigorous and fair assessment, we eliminate these confounding factors by ensuring that models can accurately comprehend the multi-modal context and respond based solely on the provided information. We elaborate on this design choice below: 1) Neutralizing Knowledge Bias: By ensuring models can derive answers directly from the context, we decouple modality preference from the interference of varying internal knowledge bases. 2) Establishing a Level Playing Field: By verifying that all models can comprehend the multi-modal context, we establish a “common ground”—a baseline of competence that ensures a fair comparison across MLLMs with diverse capabilities. Finally, we construct modality-conflict pairs to evaluate and compare the modality preferences of different MLLMs in a controlled environment.
Figure 2: Results of modality preference across different MLLMs. Left: Quantified scores for Vision, Others and Text modalities using $S_{\text{vision}}$, $S_{\text{others}}$ and $S_{\text{text}}$ as well as Vision Ratio. Right: Trends of Vision Ratio with respect to model parameter size.
@@ 3.2 Semi-automated Data Construction Pipeline
In this section, we introduce our semi-automated data construction pipeline, which follows a meticulous process to ensure the reliability of the dataset. The data is derived from the TDIUC (Kafle and Kanan, 2017), sourced from MS-COCO (Lin et al., 2014). Therefore, the data is widely adopted in model development, ensuring that most MLLMs can recognize the images. Firstly, we select the image as vision context $c^{v}$, question $q$, and answer $A^{v}$ based on the image and the caption capcap for each sample from TDIUC. Then the pipeline follows these steps:
Textual Context Construction. Given a sample including $c^{v}$, $q$, $A^{v}$ and capcap, we construct candidate contrastive textual contexts $c^{t}$ that conflict with $c^{v}$ specifically in relation to $q$ but are aligned with $c^{v}$ and capcap in terms of overall scene semantics. We prompt DeepSeekV3 (Liu et al., 2024a) and ChatGPT4o-mini (Hurst et al., 2024b) to generate a distractor answer $A^{t}$ to $q$, together with $c^{t}$ that plausibly supports $A^{t}$, using carefully crafted instructions. For each model, we generate two pairs of $A^{t}$ and $c^{t}$ to facilitate downstream data selection. To ensure that all evaluated MLLMs can correctly recognize both visual and textual contexts, we employ several weak MLLMs, such as LLaVA1.5-7B (Liu et al., 2024b) and QwenVL-7B (Bai et al., 2023), as judges to select samples that can be correctly understood with respect to $c^{v}$ and $c^{t}$.
@@
Iterative Refinement. The dataset undergoes multiple rounds of refinement through a feedback loop between textual context generation and human verification, which rectifies potential errors, thus enhancing the data quality.
MC2. To this end, we develop MC2, a modality-conflict dataset comprising 2,000 samples spanning diverse tasks. The detailed construction procedures for textual context generation, the manual annotation, the data format and dataset statistics are provided in Appendix B.
@@ 3.3 Question Design and Evaluation Metric
Question Design. We reformulate the original questions with ChatGPT-4o-mini (Hurst et al., 2024a) into a binary-choice format. To further reduce potential position bias in multi-modal inputs, we adopt a consistent evaluation strategy, similar to Liu et al. (2024d). Concretely, for each question, we construct two versions by swapping the order of the answer choices. A model’s prediction is regarded as consistent only if it selects the same answer in both versions for a sample; otherwise, it is labeled as inconsistent. Such inconsistent samples are discarded from the subsequent measurement of modality preference.
Evaluation Metric. Inspired by prior work on evaluating stylistic or knowledge-related preferences of LLMs and MLLMs through conflict-pair contexts (Li et al., 2024b; Xie et al., 2023; Liu et al., 2025), we extend this to evaluate modality preference by designing a metric that captures how MLLMs respond to conflicting signals from different modalities. More importantly, through the careful design of the benchmark, we establish as a basis that the model can reliably understand both modalities in isolation. As shown in Table 16 and 17 in Appendix, all MLLMs achieve over 95% accuracy, provided with either textual or visual context.
Building on this, our metric evaluates modality preference by assessing how the model’s responses align with textual or visual input when the two provide conflicting signals. The model’s response is then categorized based on which modality it aligns with: 1) Vision: the response aligns with the visual context; 2) Text: the response aligns with the textual context; 3) Others: the response is uncertain, or inconsistent with either modality, which is discarded from further analysis. Then, we naturally define the Vision Ratio to quantify the model’s preference toward the vision modality, defined as: ${S_{\text{vision}}}/({S_{\text{vision}}+S_{\text{text}}})$, where $S_{\text{vision}}$ or $S_{\text{text}}$ denotes the score of the vision or text modality, computed as the proportion of samples whose responses are categorized as Vision or Text across the dataset. $S_{\text{others}}$ is the proportion of samples whose responses are categorized as Others. Vision Ratio greater than 0.5 indicates that the model tends to favor visual context over text.
Figure 3: Analysis of modality preferences. Left: Trends of Vision Ratio and Attention Ratio. Middle: Vision Ratio with Respect to the proportion of multi-modal conflict-context training data and different MLLMs. Right: Relationship between visual understanding ability (quantified as the average accuracy across seven widely used benchmarks) and the modality preference (measured by Vision Ratio).
@@ 4 Modality Preferences in MLLMs
This section presents a systematic investigation into modality preference in MLLMs, structured around four key research questions: 1) Which modality do MLLMs prioritize? 2) What factors drive these preferences? 3) Can the Vision Ratio provide guidance for downstream task performance? 4) Can modality preference be controlled? This investigation helps uncover the underlying mechanisms of modality preference and leverage these insights to downstream tasks.